Hi, I'm currently on binomials. What should I do to solve this?
"When (3a+x2)6 is expanded, the coefficient of the term in x4 is 19440. Find the value of a."

Found 2 solutions by ewatrrr, Vladdroid:
Answer by ewatrrr(24785)

729x^6 + 1458a^5x^2 + 1215a^4x^4 ....
19440/1215 = 16 ⇒ a^4 = 16, a = 2

Answer by Vladdroid(91)

729x^6 + 1458a^5x^2 + 1215a^4x^4 ....
19440/1215 = 16 ⇒ a^4 = 16, a = 2